Please help me this question
y=13-4x
3x+2y=16
What is substitution

Answer by addingup(3677)

y = 13-4x this is the value of y. We will substitute in the next equation. First, set the next equation in the same format, subtract 3x from both sides:
3x+2y = 16
2y = 16-3x now in this equation substitute for y:
2(13-4x) = 16-3x Distribute the 2 over the terms inside parenthesis (Distributive Properties of multiplication):
26-8x = 16-3x Subtract 16 and add 8x on both sides:
26-16-8x+8x = 16-16-3x+8x
10 = 5x flip the equation, we want the variable on the left (it doesn't change anything):
5x = 10 divide both sides by 5
x = 2
Now solve for y:
y = 13-4x
y = 13-4(2)
y = 13-8
y = 5
------------------
Check:
3x+2y = 16
3(2)+2(5) = 16
6+10 = 16 Correct